OpenStudy (anonymous):

If a stone is tossed from the top of a 310 meter building, the height of the stone as a function of time is given by h(t) = -9.8t^2 – 10t + 310, where t is in seconds, and height is in meters. After how many seconds will the stone hit the ground? Round to the nearest hundredth’s place; include units in your answer.

OpenStudy (zarkon):

are you sure that is the function to be used?

OpenStudy (amistre64):

- 10 would indicate being tossed up first

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

or thrwon down?

OpenStudy (anonymous):

310=-9.8t^2-10t+310 =>-9.8t^2-10t=0 =>t=0 or t=-10/9.8 .......

OpenStudy (zarkon):

then this is not happening on earth :)

OpenStudy (amistre64):

-4.9 ....

OpenStudy (zarkon):

correct

OpenStudy (anonymous):

why ans is negative?

OpenStudy (zarkon):

solve h(t)=0 for t

OpenStudy (anonymous):

i think -(time) means before 4.9 sec when ball is thrown...ball is assumed to be on the ground

OpenStudy (amistre64):

the equation used for any problem dealing with gravity (on earth) comes from the fact that acceleration due to gravity is a constant. -9.8 meters per second per second. a(t) = -9.8 v(t) = -9.8 t + C, where at time 0, we have an initial velocity of -10; t=0, then C = -10 v(t) = -9.8 t -10 h(t) = -4.9 t^2 -10t + C; where at t=0, the initial height is off a building that is 310 meters tall: C = 310 h(t) = -4.9 t^2 -10t + 310

OpenStudy (amistre64):

the equation as presented in the posting, has gravity having way to large of an effect

OpenStudy (anonymous):

thanks amistre

OpenStudy (amistre64):

yep ;) at any rate, the equation should be solved for h(t) = 0 so you know at what time it takes for it to hit the ground